1. The Queen Mary is a large cruise ship that has 8,123,700 cubic feet of cargo space. Express this number in scientific notation.
answer I have-8.1237x10 to the 7th
to convert this i moved the decimal left or right until it became a number.
I got this question wrong I was told it was to the the 5th power. Is this right?
2. Georgina is collecting data on the growth of E-coli bacteria. Each E-coli cell reproduces by dividing in half every 20 minustes.
part a. If Georgina starts with 1 e-coli, how many cells will be in the sample after 120 minutes?
answer-120/120=6
6x2=18
part b. After how many minutes will there be 256 cells in sample?
answer- 256/2=7128
128x20=2560
I also got this one wrong. Not sure what I missed. These are word problems from extended responses I have every 5 weeks. These are not from my textbook. If you please help. I can redo and bring up grade. Thank you

Answer by stanbon(48558)

1. The Queen Mary is a large cruise ship that has 8,123,700 cubic feet of cargo space. Express this number in scientific notation.
answer I have-8.1237x10 to the 7th
to convert this i moved the decimal left or right until it became a number.
I got this question wrong I was told it was to the the 5th power. Is this right?
I believe it should be 8.1237 x 10^6
The "6" will move the decimal 6 to the right and
give you your original number.
-------------------------------
2. Georgina is collecting data on the growth of E-coli bacteria. Each E-coli cell reproduces by dividing in half every 20 minutes.
part a. If Georgina starts with 1 e-coli, how many cells will be in the sample after 120 minutes?
After 20 minutes there will be 2 cells
After 2*20 minutes there will be 4=2^2
After 3*20 minutes there will be 8=2^3
120/20=6
2^6=64 E-coli after 120 minutes
-----------------
part b. After how many minutes will there be 256 cells in sample?
256=2^8
8*20=160 minutes
After 160 minutes there will be 256 cells of E-coli.
-----------------
Cheers,
Stan H.